# Do PAC-Learners Learn the Marginal Distribution?

### v2

@@ Abstract
The Fundamental Theorem of PAC Learning asserts that learnability of a concept class $\mathcal{H}$ is equivalent to the uniform convergence of empirical error in $\mathcal{H}$ to its mean, or equivalently, to the problem of density estimation, learnability of the underlying marginal distribution with respect to events in $\mathcal{H}$. This seminal equivalence relies strongly on PAC learning’s ‘distribution-free’ assumption, that the adversary may choose any marginal distribution over data. Unfortunately, the distribution-free model is known to be overly adversarial in practice, failing to predict the success of modern machine learning algorithms, but without the Fundamental Theorem our theoretical understanding of learning under distributional constraints remains highly limited.
In this work, we revisit the connection between PAC learning, uniform convergence, and density estimation beyond the distribution-free setting when the adversary is restricted to choosing a marginal distribution from a known family $\mathscr{P}$. We prove that while the traditional Fundamental Theorem indeed fails, a finer-grained connection between the three fundamental notions continues to hold:
@@
    PAC-Learning is strictly sandwiched between two refined models of density estimation, both equivalent to standard density estimation in the distribution-free case, differing only in whether the learner knows the set of well-estimated events in $\mathcal{H}$.
@@
    Under reasonable assumptions on $\mathcal{H}$ and $\mathscr{P}$, density estimation is equivalent to uniform estimation, a relaxation of uniform convergence allowing non-empirical estimators.
@@ 1 Introduction
Valiant [1] and Vapnik and Chervonenkis’ [2] Probably Approximately Correct (PAC)-Learning is the foundation of modern learning theory. Traditionally, the model captures learnability of a hypothesis class $(X,H)$, a set $X$ of ‘raw data’ and a family of binary classifiers $H$, against a worst-case or ‘distribution-free’ adversary, meaning algorithmic success is always measured against the worst possible marginal distribution over data. This setting gives rise to an elegant theory of learning centered around the ‘Fundamental Theorem of PAC Learning’, stating the complexity of learning $H$ is controlled by uniform convergence, the extent to which the empirical mean of a large enough sample uniformly estimates the error of every $h\in H$ [2, 3, 4].
Unfortunately, in practice, distribution-free PAC-learning is overly adversarial. Traditional lower bounds rely on contrived data distributions not seen in practice, and modern architectures like deep nets learn much faster than PAC theory predicts [5, 6, 7]. At a theoretical level, this raises a natural question:
@@
We study a foundational variant of this question called the distribution-family model, which restricts the adversary to choosing data from a known family $\mathscr{P}$ of marginal distributions. Characterizing learnability in this setting is an old and challenging problem [2, 8, 9, 10, 11], dating back to work of Benedek and Itai [9] observing the failure of uniform convergence in the setting. Since that time, progress on the distribution family model has been largely negative, with Dudley, Kulkarni, Richardson, and Zeitouni [11] falsifying a conjectured characterization of [9] based on metric entropy, and recent work of Lechner and Ben-David [12] showing the model has no scale-insensitive combinatorial characterization.
In this work, we revisit the conventional knowledge that the Fundamental Theorem fails beyond the distribution-free setting and aim to offer a more nuanced view of the connection between uniform convergence and its variants and learning under distributional assumptions. Toward this end, it is constructive to consider a rephrasing of the Fundamental Theorem in terms of a standard statistical framework: density estimation. Given a triple $(\mathscr{P},X,H)$ and oracle access to a distribution $D\in\mathscr{P}$, the density estimation problem asks the learner to output $D^{\prime}\in\mathscr{P}$ that is close to $D$ in the total variation metric with respect to $H$: [^1] Here $H\Delta H$ denotes the set of symmetric differences in $H$.
@@
This is natural in the context of supervised learning where events in $H\Delta H$ measure the error between hypotheses, and is clearly closely related to uniform convergence. Indeed, the following folklore rephrasing [13, 14, 15, 16] of the Fundamental Theorem states that density estimation, uniform convergence, and PAC learning are all equivalent in the distribution-free model.
@@
While Benedek and Itai [9] observe the latter equivalence breaks in the distribution-family model, their example does not preclude an equivalence with density estimation, which is a common strategy in learning with distributional assumptions (see e.g. [17]). We will work with the following weakening (equivalent in the distribution free setting), which is clearly sufficient to learn so long as each $(D,X,H)$ in $(\mathscr{P},X,H)$ is individually learnable. [^2] Formally here we mean the $(D,X,H)$ are learnable in a uniformly bounded number of samples, which Benedek and Itai [9] proved is equivalent to uniformly bounded metric entropy. If this does not hold $(\mathscr{P},X,H)$ is trivially unlearnable, and the problem becomes uninteresting. Throughout this work, we will always assume this condition.
@@
$(\mathscr{P},X,H)$ has an intermediate density estimator (IDE) if there exists a learner and (sample dependent) subset $G\subset H$ satisfying:
@@
Unlike our previous refinement, it is no longer apriori clear how weak density estimation relates to PAC Learning, even in the distribution-free case. The notion is clearly necessary, but without knowledge of $G$, it is not obvious how to use such an estimator to construct a learner. Nevertheless, we show weak density estimation case requires finite VC dimension in the distribution-free (Proposition 4.1), and therefore remains equivalent to PAC Learning.
Returning to the distribution-family model, we show that unlike intermediate estimation, weak estimation is necessary to PAC-Learn (Proposition 4.2). However, unlike the distribution-free case, we exhibit a class that is weakly estimatable but cannot be PAC Learned, breaking the equivalence in the other direction.
@@
Taken together, these results lead to our main theorem: PAC-Learning with distributional assumptions is strictly sandwiched between two close refinements of density estimation, both equivalent to standard estimation and learning in the distribution-free setting.
@@
The following (in)equivalences hold:
  1. 1.
    In the distribution-free setting
    Weak Density Estimation $=$ PAC-Learning $=$ Intermediate Density Estimation
  2. 2.
@@
    Weak Density Estimation $\subsetneq$ PAC-Learning $\subsetneq$ Intermediate Density Estimation.
Finally, we return to the topic of uniform convergence and its relation to density estimation and learning. While uniform convergence is strictly stronger than learning and density estimation in the distribution family model, a natural relaxation called uniform estimation allowing for non-empirical estimators is more closely related. We prove that when $(\mathscr{P},X,H)$ has uniformly bounded entropy, the two notions are equivalent.
@@
We remark that while this connection is standard in the distribution-free setting [13], the argument in the distribution-family model is more subtle and combines standard methods in density estimation [13, 14] with recent connections between covering and learning with distributional assumptions [18]. We also note Proposition 1.7 is false for the natural analog of ‘weak estimation’ that requires estimating error for most hypotheses in the class without knowing the set of ‘good’ estimates. Such a guarantee is immediate from Chernoff, while weak density estimation requires finite VC dimension in the distribution-free case.
Taken together, Theorem 1.6 and Proposition 1.7 provide a finer-grained picture of the classical connection between supervised learning, uniform convergence, and density estimation. It remains an interesting open problem to prove further positive results for learning with distributional assumptions as well. One promising direction is to ask whether there are natural assumptions on the class $\mathscr{P}$ over which uniform estimation and density estimation are necessary. Another interesting direction is to explore classical paradigms such as sample compression or stability, known to characterize learnability in other settings where uniform convergence fails [19, 20].
Roadmap:
In Section 2 we cover preliminaries and related work, in Section 3 we study intermediate density estimation, in Section 4 we study weak density estimation, and in Section 5 we relate density estimation to other classical notions in learning theory including uniform convergence, uniform estimation, and the distribution-free setting.
@@ 2 Background and Related Work
2.1 PAC-Learning
We recall some standard background in PAC-learning, starting with the definition in the distribution-family setting. We study learnability of triples $(\mathscr{P},X,H)$ where $X$ is a set, $H$ is a family of binary classifiers over $X$, and $\mathscr{P}$ is a family of distributions over $X$.
@@ 2.1 PAC-Learning
where $d_{D}(\cdot,\cdot)$ is the classification distance
$d_{D}(h,h^{\prime})=\Pr_{x\sim D}[h(x)\neq h^{\prime}(x)].$
This variant is often called the realizable setting, since there must exist a function in our class $H$ that realizes the adversary’s choice exactly. While there is a well-studied model called the agnostic setting that relaxes this assumption, it is known to be statistically equivalent to the realizable case [18], so we focus on this setting throughout.
2.2 Uniform Convergence
Characterizations of the distribution-family model are known in several special cases. Most famously, the ‘distribution-free’ setting where $\mathscr{P}$ consists of the set of all distributions is characterized [^3] Under weak measurability assumptions, see e.g. [3]. by a combinatorial parameter called VC-dimension [2, 3, 4], or equivalently by a property called uniform convergence.
@@ 2.2 Uniform Convergence
In this work, we also consider a relaxed form of uniform convergence that allows for use of estimators beyond the empirical error.
Definition 2.3 (Uniform Estimation).
We say $(\mathscr{P},X,H)$ satisfies the uniform estimation property if $\forall\varepsilon,\delta>0$ there exists a family of estimators $\{\mathcal{E}_{S}\}_{S\in(X\times\{0,1\})^{*}}:H\to\mathbb{R}_{+}$ and a function $n=n_{\text{UE}}(\varepsilon,\delta)$ such that for all $D\in\mathscr{P}$, and $h\in H$, the estimated error of every $h^{\prime}\in H$ approaches its true error uniformly:
We call such a family an $(\varepsilon,\delta)$-uniform estimator.
Bounds on uniform convergence or estimation in the distribution-free case typically depend on a complexity measure of the class called the growth function.
Definition 2.4 (Growth Function).
Given a class $(X,H)$, the growth function $\Pi_{H}(n)$ is the maximum number of labelings attained by $H$ on any subset of $X$ of size $n$:
$\Pi_{H}(n)\coloneqq\max_{S\subset X:|S|=n}|\{H|_{S}\}|$
where $H|_{S}$ is the restriction of $H$ to $S$.
Finally, we note that in the distribution-free setting the growth function (and therefore uniform convergence) is controlled by a combinatorial parameter of the class called the VC dimension.
Definition 2.5 (VC Dimension).
The VC dimension of a class $(X,H)$ is the maximum cardinality of any set $S=\{x_{1},\ldots,x_{d}\}$ in $X$ such that all possible labelings of $S$ are realizable by $H$.
2.3 Metric Entropy
Unlike the distribution-free setting, uniform convergence fails to characterize the fixed-distribution model. Benedek and Itai [9] gave a characterization in this setting based on finite covers.
@@ 2.3 Metric Entropy
$\forall h\in H,\exists c\in C\text{:}\ \ d_{D}(c,h)\leq\varepsilon.$
@@
Benedek and Itai observed that any class $(D,X,H)$ is PAC-learnable if and only if it has finite metric entropy. This lead them to conjecture that $(\mathscr{P},X,H)$ is learnable if and only if it has uniformly bounded metric entropy across $\mathscr{P}$.
@@
We say $(\mathscr{P},X,H)$ has uniformly bounded metric entropy (UBME) if there exists a function $m_{\text{UB}}(\varepsilon)$ such that for all $\varepsilon>0$ and $D\in\mathscr{P}$, $(D,X,H)$ has an $\varepsilon$-cover of size at most $m_{\text{UB}}(\varepsilon)$. We call minimum such $m_{\text{UB}}$ the metric entropy of $(\mathscr{P},X,H)$.
While UBME is known to be equivalent to VC dimension in the distribution-free setting [4], Benedek and Itai’s conjecture was disproved in the general distribution family setting by Dudley, Kulkarni, Richardson, and Zeitouni (DKRZ) [11], who exhibited an unlearnable class with UBME.
@@ 2.4 Density Estimation
Our work considers the connections between the distribution-family model and several new variants of density estimation. We first recall the standard model adapted to our setting where the learner’s goal is simply to output an estimate to the adversary’s choice of $D\in\mathscr{P}$ with respect to the $H\Delta H$-distance (recall $H\Delta H=\{h\Delta h^{\prime}:h,h^{\prime}\in H\}$ is the set of symmetric differences of pairs of hypotheses in $H$).
@@
In the context of PAC-learning, standard density estimation (even restricted to $H\Delta H$) is still needlessly strong. In particular, the learner really need only estimate the distances for most $h\in H$, provided it knows the set of well-estimated events. We call such a guarantee intermediate density estimation (IDE).
@@
While a weakening of the standard model, requiring knowledge of the well-estimated events is still a strong assumption (and one we exploit in Section 3 to construct learnable classes failing IDE). With this in mind, we introduce the notion of weak density estimation (WDE) which drops this constraint.
@@
Finally in the analysis of density estimation we will frequently use to the empirical distance estimator corresponding to an unlabeled sample, defined as:
Definition 2.12 (Empirical Distance Estimator).
Given an unlabeled sample $T\in X^{*}$ and hypothesis class $(X,H)$, the empirical distance estimator is
$\hat{d}_{T}(h,h^{\prime})\coloneqq\frac{1}{|T|}\sum\limits_{x\in T}\mathbf{1}\{h(x)\neq h^{\prime}(x)\}.$
@@ 2.5 Further Related Work
The distribution-family model can be seen as implicit in the seminal works of Vapnik and Chervonenkis [8, 2], who studied general models of learning under convergence of empirical means. These bounds were later sharpened by Natarajan [10]. In 1991, Benedek and Itai [9] explicitly introduced the distribution-family model, and conjectured a characterization based on metric entropy which was falsified by Dudley, Kulkarni, Richardson, and Zeitouni [11] several years later. Finally, in 1997 Kulkarni and Vidyasagar [21] observed the sufficiency of UBME in the finite and distribution-free settings could be slightly generalized to finitely coverable families and families with an interior point (roughly these are locally distribution-free).
Recently, Lechner and Ben-David [12] proved the distribution-family model has no scale-insensitive characterizing dimension. We note that an earlier version of this manuscript appeared publicly before the work of Lechner and Ben-David, and has been updated to reflect their result. Finally, while not strictly related, it is also worth mentioning the recent universal learning model [22], which may also be viewed as restricting the adversary in a way that escapes traditional VC lower bounds.
Density Estimation.
@@
Density estimation (also called ‘improper’ distribution learning [23]) is among the most classical problem in statistics and a special case of Vapnik’s general learning setting [24]. We do not attempt to give a full summary here, and refer the reader to the book of Lugosi and Devroye [14] for the most relevant overview. Our techniques draw from a line of work using combinatorial methods in density estimation originating in the seminal paper of Yatracos [13], further developed by Lugosi and Devroye [14] and the more recent distribution learning literature (see e.g. [15, 23, 16]). Several of our results leverage classical ideas from this line, in particular the seminal idea of Yatracos lifting empirical estimates on covers to density estimates on the entire space.
@@
The relation between variants of density estimation and supervised learning is also studied in [25] under the ‘Known-Labeling Classifier Learning’ (KLCL) model. This setting, which looks at distribution-free learning when a (non-realizable) labeling is known to the learner, is incomparable to our own and requires a different set of techniques.
Unsupervised learning.
@@
Our work also takes inspiration from unsupervised techniques in the machine learning literature. Especially relevant to our setting are multi-stage supervised classifiers that infer information from unlabeled data before applying traditional supervised approaches to achieve state-of-the-art accuracy and sample guarantees [26, 27, 28, 29, 30, 17]. Our approach is perhaps most explicitly inspired by techniques of Balcan and Blum [26] in semi-supervised learning, but is also related to similar ideas used in transfer learning [31], active learning [32, 33, 34, 35, 36, 37, 29], privacy [27, 28], and even distribution-free learning of halfspaces [38, 39, 40, 41, 30]. Most recently, a similar approach was used by Hopkins, Kane, Lovett, and Mahajan [18] to give generic reductions from agnostic to realizable settings across many variants of supervised learning (including the distribution-family model). Our notion of weak density estimation draws inspiration from their work, which focuses on semi-supervised techniques without uniform convergence.
Characterizations Beyond Uniform Convergence

### v3

@@ Abstract
The Fundamental Theorem of PAC Learning asserts that learnability of a concept class $H$ is equivalent to the uniform convergence of empirical error in $H$ to its mean, or equivalently, to the problem of density estimation, learnability of the underlying marginal distribution with respect to events in $H$. This seminal equivalence relies strongly on PAC learning’s ‘distribution-free’ assumption, that the adversary may choose any marginal distribution over data. Unfortunately, the distribution-free model is known to be overly adversarial in practice, failing to predict the success of modern machine learning algorithms, but without the Fundamental Theorem our theoretical understanding of learning under distributional constraints remains highly limited.
In this work, we revisit the connection between PAC learning, uniform convergence, and density estimation beyond the distribution-free setting when the adversary is restricted to choosing a marginal distribution from a known family $\mathscr{P}$. We prove that while the traditional Fundamental Theorem fails, a finer-grained connection between the three fundamental notions continues to hold:
@@
    PAC-Learning is strictly sandwiched between two relaxed models of density estimation, differing only in whether the learner knows the set of well-estimated events in $\mathcal{H}$.
@@
    Under reasonable assumptions on $H$ and $\mathscr{P}$, density estimation is equivalent to uniform estimation, a weakening of uniform convergence allowing non-empirical estimators.
@@
keywords:
PAC-Learning, Density Estimation, Uniform Convergence
@@ 1 Introduction
[Valiant (1984)] and [Vapnik and Chervonenkis (1974)] Probably Approximately Correct (PAC)-Learning is the foundation of modern learning theory. Traditionally, the model captures learnability of a hypothesis class $(X,H)$, a set $X$ of ‘raw data’ and a family of binary classifiers $H$, against a worst-case or ‘distribution-free’ adversary, meaning algorithmic success is always measured against the worst possible marginal distribution over data. This setting gives rise to an elegant theory of learning centered around the ‘Fundamental Theorem of PAC Learning’, stating the complexity of learning $H$ is controlled by uniform convergence, the extent to which the empirical mean of a large enough sample uniformly estimates the error of every $h\in H$ (Vapnik and Chervonenkis, 1974; Blumer et al., 1989; Haussler, 1992).
Unfortunately, in practice, distribution-free PAC-learning is overly adversarial. Traditional lower bounds rely on contrived data distributions not seen in practice, and modern architectures like deep nets learn much faster than PAC theory predicts (Neyshabur et al., 2014; Zhang et al., 2021; Nagarajan and Kolter, 2019). At a theoretical level, this raises a natural question:
@@
We study a foundational variant of this question called the distribution-family model, which restricts the adversary to choosing data from a known family $\mathscr{P}$ of marginal distributions. Characterizing learnability in this setting is an old and challenging problem (Vapnik and Chervonenkis, 1974, 1971; Benedek and Itai, 1991; Natarajan, 1992; Dudley et al., 1994), dating back to work of Benedek and Itai (1991) observing the failure of uniform convergence in the setting. Since that time, progress on the distribution family model has been largely negative, with Dudley et al. (1994) falsifying a conjectured characterization of Benedek and Itai (1991) based on metric entropy, and recent work of Lechner et al. (2023) showing the model has no scale-insensitive combinatorial characterization.
In this work, we revisit the conventional knowledge that the Fundamental Theorem fails beyond the distribution-free setting and aim to offer a more nuanced view of the connection between uniform convergence and its variants and learning under distributional assumptions. Toward this end, it is constructive to consider a rephrasing of the Fundamental Theorem in terms of a standard statistical framework: density estimation. Given a triple $(\mathscr{P},X,H)$ and oracle access to a distribution $D\in\mathscr{P}$, the density estimation problem asks the learner to output $D^{\prime}\in\mathscr{P}$ that is close to $D$ in the total variation metric with respect to $H$: [^1] Here $H\Delta H\coloneqq\{h\Delta h^{\prime}:h,h^{\prime}\in H\}$ denotes the set of symmetric differences in $H$.
@@
This is natural in the context of supervised learning where events in $H\Delta H$ measure the error between hypotheses, and is clearly closely related to uniform convergence. Indeed, the following folklore rephrasing (Yatracos, 1985; Devroye and Lugosi, 2001; Chan et al., 2014; Ashtiani et al., 2018) of the Fundamental Theorem states that density estimation, uniform convergence, and PAC learning are all equivalent in the distribution-free model.
@@
While Benedek and Itai (1991) observe the latter equivalence breaks in the distribution-family model, their example does not contradict a potential equivalence with density estimation. Indeed, density estimation is a common strategy in learning with distributional assumptions (see e.g. Blanc et al. (2023)). We will work with the following weakening (equivalent in the distribution free setting), which is clearly sufficient to learn so long as each $(D,X,H)$ in $(\mathscr{P},X,H)$ is individually learnable. [^2] Formally here we mean the $(D,X,H)$ are learnable in a uniformly bounded number of samples, which Benedek and Itai (1991) proved is equivalent to $(\mathscr{P},X,H)$ having uniformly bounded metric entropy (see Definition 2.5). If this does not hold $(\mathscr{P},X,H)$ is trivially unlearnable, and the problem becomes uninteresting. Throughout this work, we will always assume this condition.
@@
$(\mathscr{P},X,H)$ has an intermediate density estimator (IDE) if there exist a learner and (sample dependent) subset $G\subset H$ satisfying:
@@
Unlike the intermediate model, we prove weak estimation is necessary to PAC-learn (see Proposition 3.9). One might then hope the condition also remains sufficient, but this turns out to be false in general as well.
@@
Taken together, these results lead to our main theorem: PAC-Learning with distributional assumptions is strictly sandwiched between two close models of density estimation.
@@
[…]
Finally, we return to the topic of uniform convergence and its relation to density estimation and learning. While uniform convergence is strictly stronger than learning and density estimation in the distribution family model, a natural relaxation called uniform estimation which allows for non empirical estimators is more closely related. Indeed, we prove that when $(\mathscr{P},X,H)$ has uniformly bounded metric entropy, the two notions are equivalent.
@@
We emphasize Proposition 1.7 only holds under the assumption that each $(D,X,H)$ in the class is learnable in a uniformly bounded number of samples. Indeed the statement is trivially false otherwise, see Section 6 for further discussion.
We remark that while this connection is standard in the distribution-free setting (Yatracos, 1985), the argument in the distribution-family model is more subtle and combines standard methods in density estimation (Yatracos, 1985; Devroye and Lugosi, 2001) with recent connections between covering and learning with distributional assumptions (Hopkins et al., 2024). We conjecture that Proposition 1.7 should actually fail for the corresponding ‘weak’ estimation variant. Here weak uniform estimation (estimating error for most hypotheses in the class without knowing the set of ‘good’ estimates) is immediate from Chernoff. On the other hand, weak density estimation asks not just for a good estimate of most $d_{D}(h,h^{\prime})$, but rather for a large set $G$ where every distance $d_{D}(g,g^{\prime})$ is well estimated. This is a more involved condition, and we leave a characterization of classes satisfying the property as an open problem for future work.
Taken together, Theorem 1.6 and Proposition 1.7 provide a finer-grained picture of the classical connection between supervised learning, uniform convergence, and density estimation. It remains an interesting open problem to prove further positive results for learning with distributional assumptions as well. One promising direction is to ask whether there are natural assumptions on the class $\mathscr{P}$ over which uniform estimation and density estimation are necessary. Another interesting direction is to explore classical paradigms such as sample compression or stability, known to characterize learnability in other settings where uniform convergence fails (Shalev-Shwartz et al., 2010; David et al., 2016).
@@ 2 Background and Related Work
2.1 PAC Learning, Uniform Convergence, and Metric Entropy
Let $X$ be a set, $H$ a family of binary classifiers over $X$, and $\mathscr{P}$ a family of distributions over $X$.
@@ 2.1 PAC Learning, Uniform Convergence, and Metric Entropy
where $d_{D}(h,h^{\prime})=\Pr_{x\sim D}[h(x)\neq h^{\prime}(x)]$ is the classification distance.
In general one might also consider an agnostic model allowing arbitrary labelings, but this is statistically equivalent to the above (Hopkins et al., 2024) so we focus on the simpler realizable case.
Characterizations of the distribution-family model are known in several special cases. Most famously, the ‘distribution-free’ setting where $\mathscr{P}$ consists of the set of all distributions is characterized by uniform convergence (Vapnik and Chervonenkis, 1974; Blumer et al., 1989).
@@
Bounds on uniform convergence or estimation in the distribution-free case typically depend on a complexity measure of the class called the growth function $\Pi_{H}(n)\coloneqq\max_{S\subset X:|S|=n}|\{H|_{S}\}|$, where $H|_{S}$ is the family of binary classifiers on the sample $S$ realized by $H$.
When $\mathscr{P}$ consists of just a single distribution, uniform convergence fails to characterize learnability. Here Benedek and Itai (1991) instead give a characterization based on finite covers.
@@
$\forall h\in H,\exists c\in C\text{:}\ \ d_{D}(c,h)\leq\varepsilon$
@@
Benedek and Itai (1991) proved $(D,X,H)$ is PAC-learnable if and only if it has finite metric entropy, and conjectured a similar equivalence for the following generalization to $(\mathscr{P},X,H)$.
@@
We say $(\mathscr{P},X,H)$ has uniformly bounded metric entropy (UBME) if there exists $m_{\text{UB}}(\varepsilon)$ such that for all $\varepsilon>0$ and $D\in\mathscr{P}$, $(D,X,H)$ has an $\varepsilon$-cover of size at most $m_{\text{UB}}(\varepsilon)$. We call minimum such $m_{\text{UB}}$ the metric entropy of $(\mathscr{P},X,H)$.
UBME is known to be equivalent to PAC-learnability in the distribution-free setting (Haussler, 1992), but is not sufficient in the general distribution-family setting (Dudley et al., 1994).
@@ 2.2 Density Estimation
Our work considers connections between the distribution-family model and several new variants of density estimation. We first recall the standard model adapted to our setting where the learner wishes to output an estimate to the adversary’s choice of $D\in\mathscr{P}$ with respect to the $H\Delta H$-distance.
@@
In the context of PAC-learning, standard density estimation (even restricted to $H\Delta H$) is needlessly strong: the learner really need only estimate the distances for most $h\in H$, provided it knows the set of well-estimated events. We call this condition intermediate density estimation.
@@
While a weakening of the standard model, requiring knowledge of well-estimated events is still a strong assumption (and one we exploit in Section 3.1 to construct learnable classes failing IDE). With this in mind, we introduce weak density estimation (WDE) which drops this constraint.
@@ 2.3 Further Related Work
Distribution Family Model. The distribution-family model is implicit in the seminal works of Vapnik and Chervonenkis (1971, 1974), who studied general models of learning under convergence of empirical means. These bounds were later sharpened by Natarajan (1992). In 1991, Benedek and Itai (1991) explicitly introduced the model and conjectured its characterization by UBME which was falsified by Dudley et al. (1994) several years later. In 1997 Kulkarni and Vidyasagar (1997) observed the sufficiency of UBME in the finite and distribution-free settings could be slightly generalized to finitely coverable families and families with an interior point. Recently, [^3] A version of this work appeared publicly before Lechner et al. (2023) and has been updated to reflect their result. Lechner et al. (2023) proved the distribution-family model has no scale-insensitive characterizing dimension.
@@
Density Estimation. Density estimation (also called ‘improper’ distribution learning (Diakonikolas, 2016)) is among the most classical problems in statistics. We refer the reader to Devroye and Lugosi (2001) for the most relevant overview of the area. Our techniques draw from combinatorial methods in density estimation due to Yatracos (1985) further developed in Devroye and Lugosi (2001) and the distribution learning literature (Chan et al., 2014; Diakonikolas, 2016; Ashtiani et al., 2018). Several of our results leverage ideas from this line, in particular Yatracos’ method of lifting empirical estimates on covers to full density estimates.
@@
The relation between variants of density estimation and supervised learning is also studied in Ben-David and Ben-David (2011) under the ‘Known-Labeling Classifier Learning’ (KLCL) model. This setting, which looks at distribution-free learning when a (non-realizable) labeling is known to the learner, is incomparable to our own and requires a different set of techniques.
@@
Unsupervised learning. Our work also takes inspiration from unsupervised techniques in machine learning, especially multi-stage classifiers that infer information from unlabeled data before applying supervised methods (Balcan and Blum, 2010; Beimel et al., 2013; Alon et al., 2019; Hopkins et al., 2020c; Diakonikolas et al., 2022; Blanc et al., 2023), but also by similar techniques from transfer learning (Pan and Yang, 2009), active learning (Balcan et al., 2007; Hanneke and Yang, 2015; Kane et al., 2017; Hopkins et al., 2020a, b, c), privacy (Beimel et al., 2013; Alon et al., 2019), and even distribution-free learning of halfspaces (Diakonikolas et al., 2021, 2022). Most recently, a similar approach was used by Hopkins et al. (2024) to give generic reductions from agnostic to realizable settings across many variants of supervised learning (including the distribution-family model). Our notion of weak density estimation draws inspiration from their work.